Chemists have determined that 18.0 g of water consists of 6.02 x 10 23 molecule. Assuming that a tbsp holds 14.79 mL of water, determine the number of water molecules the tbsp can hold? (density water = 1g/mL)
1
Expert's answer
2020-09-21T06:37:40-0400

The mass of water in a table spoon is:

m(water) = d(water) × V(water)

where d - density, V - volume.

From here:

m(water) = 1 g/mL × 14.79 mL = 14.79 g

As 18 g of water consists of 6.02 × 10²³ molecules, 14.79 g of water consists of:

n(water) = 6.02 × 10²³ × 14.79 g / 18.0 g = 4.95 × 10²³


Answer: 4.95 × 10²³